OpenStudy (anonymous):

solve. 2(a-1)^2-1=31

OpenStudy (anonymous):

add one to both sides, get \[2(a-1)^2=32\] divide by 2, get \[(a-1)^2=16\] take the square root get \[a-1=4\] or \[a-1=-4\] add 1 get \[a=5\] or \[a=-3\]